# Populations of tidal and pulsating variables in eclipsing binaries

### v1

@@ Abstract
In the current age of space-based precision photometry, photometric observations provide the most complete time-domain coverage of binary star systems. Kepler (Borucki et al., 2010), TESS (Ricker et al., 2015), and the upcoming PLATO mission (Rauer et al., 2025) are designed to provide high-precision photometry for large numbers of targets, resulting in the detection and characterisation of tens of thousands of eclipsing binaries (EBs, Prša et al. 2011; Slawson & Doyle 2011; Kirk et al. 2016; Prša et al. 2022). These EBs allow invaluable constraints to be placed on the binary components and their orbital properties, information that is highly complementary with surface properties obtained through traditional spectroscopic follow-up or, where unavailable or impractical, estimates based on Gaia’s BP/RP low resolution spectra (Gaia Collaboration et al., 2023b).
@@
A natural synergy between photometric observations of EBs is that with asteroseismology, as the same high-quality time series that allows measurement of eclipse properties also captures flux variations due to pulsations (Gaia Collaboration et al., 2023a). At the level of individual stars, this synergy has been proved to be powerful; the combination of EB constraints (Hambleton et al., 2013; Schmid & Aerts, 2016; Guo et al., 2019; Sekaran et al., 2021; Guo, 2021; Aerts & Tkachenko, 2024; Kemp et al., 2024, 2025) and asteroseismic forward modelling (Aerts et al., 2018) can provide an invaluable window on fundamental stellar mixing physics (Pedersen et al., 2021; Michielsen et al., 2021, 2023; Michielsen, 2024). These and similar studies rely heavily on the sensitivity of gravity-mode (g-mode) pulsators to the near-core properties of stars with convective cores and radiative envelopes (main sequence stars with masses greater than 1.2 M⊙). Pressure-mode (p-mode) pulsators (Bowman et al., 2016; Murphy et al., 2019, 2024) offer complementary information about the outer envelope, making stars in which both g- and p-modes are detected (hybrid pulsators) particularly valuable (Kurtz et al., 2014; Mombarg et al., 2020; Audenaert & Tkachenko, 2022; Kemp et al., 2024, 2025).
@@
At the population level, model-independent measurements of the stellar rotation in the near-core region have revealed that main sequence g-mode pulsators are (near-) rigid rotators (Aerts, 2021). Further, using nothing but the dominant g-mode oscillation frequency, an estimate of the stellar mass and near-core rotation rate can be made even from sparse Gaia photometry (Hey & Aerts, 2024; Aerts et al., 2025).
@@
Recently, large numbers of eclipsing binaries have been identified and have had their eclipse properties measured using TESS photometry (IJspeert et al., 2024a) (IJ24). Previous analysis of the asteroseismic potential of this sample was limited to an automated detection procedure based on the number of significant frequencies detected that were not associated with eclipse harmonics. We aim to compile a robust, manually vetted sample of pulsating EBs that is well-characterised in terms of their orbital and stellar properties. Of particular interest is the degree of similarity between the distributions of the subpopulations of pulsating vs non-pulsating EBs, and understanding how binarity can induce bias in the inference of stellar properties. Such a well characterised sample is ideal for applying novel asteroseismic inference methodologies such as the dominant-mode-based estimator for the near-core rotation rate in Aerts et al. (2025). Further, it provides a starting point for more traditional asteroseismic analyses for targets of high interest and data quality.
@@
In this work, we perform manual vetting of the IJ24 sample in order to identify candidate EB pulsators, and use asteroseismically calibrated stellar models from Mombarg et al. (2024b) (M24) to estimate stellar properties for the sample based on their Gaia effective temperature and luminosity estimates. We explore if, and if so how, binarity can result in systematic bias to these results, and confront our mass inferences with 2MASS (Skrutskie et al., 2006) -magnitude measurements in order to better understand the efficacy of our methodology.
@@ 2.1 Sample definition
Beginning with the 69085-strong eclipsing binary sample of IJ24, we first recover IJ24’s high-quality sample of over 14377111The precise number of stars in our sample differs slightly from IJ24 as we filter both duplicate TESS and Gaia entries, while IJ24 only filters duplicate TESS entries stars. This is achieved by demanding that the automated STAR SHADOW (IJspeert et al., 2024b) eclipse analyses of IJ24 have completed successfully, filtering for duplicate TESS and Gaia entries, and removing all binaries with that are consistent with 0 eccentricity to within . System lumjinosities are calculated using Gaia DR3 parallax measurements, while other Gaia properties (most notably the effective temperatures) are drawn from esp-hs where available, as it is most suitable for hot stars (7000 K ¡ ¡ 50000 K, Gaia Collaboration et al. 2023b), and gsp-phot otherwise.
Using the frequency analysis of IJ24, we filtered for stars which have at least one candidate independent frequency of at least 4 mmag. This threshold was designed to enrich our manually vetted sample with pulsators (Gaia Collaboration et al., 2023a), as well as improve compatibility with existing empirical asteroseismic methods in the literature (Hey & Aerts, 2024; Aerts et al., 2025) relying on dominant g-mode oscillations of at least 4 mmag in amplitude. The 4 mmag limit was originally derived from Gaia DR3 time series (Gaia Collaboration et al., 2023a), and so isn’t strictly applicable to TESS photometry. For this reason, we only use this limit for the pre-selection phase resulting in our 4000-strong subsample that we subject to manual vetting; it does not form part of the manual vetting criteria. Most of the scientific discussion in this work will focus on this 4000-target subsample, but note that the stellar property inference was conducted on the entire 14377-target sample of IJ24.
@@
Our 4000-binary subsample of IJ24 was manually inspected in the time and Fourier domains for candidate independent g-mode (751), p-mode (131), or hybrid (48) pulsators. This inspection was done on one year of TESS QLP (Huang et al., 2020a, b; Kunimoto et al., 2021) or SPOC (Kunimoto et al., 2021) data per target, where the data used was selected based on the following priority ranking from highest to lowest:
@@ 2.2 Stellar parameter estimation
Using conditional normalising flows (CNFs, Winkler et al. 2019; Hon et al. 2024) trained on the M24 stellar evolution models, we estimate masses (), convective core masses (), radii (), and central H fractions () for the 12508 targets with Gaia effective temperatures and luminosities. This estimation is based on the previously described Gaia log and log estimates, and is performed for the three metallicities () available, marginalising over the initial (rigid) rotations () and values for the exponential overshoot () varied in the M24 grids.
@@
where is the height above the convective boundary, is the pressure scale height at the convective boundary, and is the diffusivity coefficient at the edge of the convective boundary (Freytag et al., 1996; Herwig, 2000; Mombarg et al., 2024a). Following M24, we uniformly resample the evolution models in to avoid bias towards the zero-age main sequence (ZAMS) and terminal-age main sequence (TAMS) caused by the numerics of the evolution models. For all targets, a kernel density estimate (KDE) is then performed to obtain the maximum likelihood estimate and its uncertainty by uniformly distributing 800 points in each stellar parameter (, , , log, , and ).
@@ 3.1 Basic stellar and orbital properties
The distribution of our stars across the Hertzsprung-Russel diagram (HRD) is shown in Fig. 1. The upper left panel also shows solar metallicity (=0.014) evolutionary tracks from 1.2 ⊙ to 9 M⊙ from M24. P-values for 2-sample Kolmogorov-Smirnov (KS) tests (Massey Jr, 1951; Hodges, 1958) applied to all subsample pairings can be found in Table 2. Unless otherwise stated, the p-values discussed in the main text always adopt the null hypothesis that the relevant pairing of subsamples is drawn from the same underlying distribution; hence, a high p-value implies a high degree of similarity between the distributions, while a low value indicates significant discrepancies. The lower panels highlight the distribution of the g-mode, p-mode, and hybrid pulsators.
@@
The distribution of the g-mode pulsators is relatively continuous across the HRD; while the 2-sample KS test probability with the non-pulsator class is low (0.045 for log), it is also significantly higher than that of either the p-mode or hybrid samples (0.006 and 0.001, respectively). Qualitatively, there is a slight bimodality in log for the g-mode stars that is absent for the non-pulsators. This may be the signature of the Dor (Guzik et al., 2000; Dupret et al., 2005; Uytterhoeven et al., 2011; Bouabid et al., 2013) and slowly-pulsating B-star (SPB, Waelkens 1991; Dziembowski et al. 1993; Gautschy & Saio 1993; Miglio et al. 2007) instability regions. However, we still must conclude that the distribution of g-mode pulsators across the HRD appears more similar to the distribution of non-pulsators than we would expect for a (pure) sample of g-mode pulsators confined to the traditional Dor, SPB, and Cep (Dziembowski & Pamiatnykh, 1993; Stankov & Handler, 2005) instability regions. Pulsators outside these regions have previously been found by combining Gaia astrometry with Kepler or TESS photometry (Gaia Collaboration et al., 2023b; Balona, 2023; Hey & Aerts, 2024; Fritzewski et al., 2024; Mombarg et al., 2024a). It is possbile that this blurring of the traditional instability regions could be due to fundamental limitations to the precision and accuracy of Gaia estimates in the SPB regime. However, it is at least as plausible that this phenomena is due to the physical limitations of our excitation models, which are missing proper treatments for physics such as stellar rotation and tides Aerts & Tkachenko (2024).
In the context of our eclipsing binary sample, a key question is whether tidal interactions (Fuller & Lai, 2012; Fuller, 2017; Guo, 2021; Aerts & Tkachenko, 2024) are a dominant force for the g-mode pulsator classes becoming blurred. Alternative explanations include non-binary related missing physics in mode excitation models, such as rotation, but also the possibility of a relatively high level of contamination by rotational variables, which, even with manual vetting, can be difficult to distinguish from genuine g-mode pulsators in TESS photometry (Audenaert et al., 2021; Hey & Aerts, 2024).
@@
The period-eccentricity distribution of the 4000-target vetted sample is shown in Fig. 2. G-mode pulsators, including hybrids, are preferentially found at short orbital periods compared to both the non-pulsator subclass and the p-mode pulsators. This contrasts with Fig. 1, which showed that in terms of stellar surface properties, the distribution of the hybrids and p-mode were far more similar than that of hybrids and g-modes. The p-mode pulsators peak at longer orbital periods and generally have a flatter distribution that is more consistent with the non-pulsator subclass (p-value = )) than either the g-mode (p-value = ) or hybrid pulsators (p-value = )). This preference for short-period, circular binaries is consistent with increased tidal efficiency in the presence of g-mode pulsations, which would lead to more rapid spin-orbit coupling and circularisation (Rogers et al., 2013; Ogilvie, 2014).
@@
Comparing the upper and lower panels of Fig. 3, there does not appear to be a strong correlation between tidal variability and the g-mode pulsators. This is not consistent with tidal effects dominantly affecting the excitation of g-modes in our sample. If this were the case, we would expect more g-mode pulsators in systems where tidal variability is the strongest. Instead, the g-mode pulsators are distributed relatively evenly between low and moderate values (0-0.4) of our tidal morphology parameter. Comparing the different pulsator classes (see Fig. 11), we note that the g-mode pulsators prefer higher values of the tidal-morphology parameter than the p-mode pulsators (p-value = 0.004), while the hybrids closely match the g-mode distribution (p-value = 0.73, compared to a p-mode — hybrid p-value of 0.03). While this does constitute evidence for g-mode pulsators being affected by tides more than targets without g-mode pulsations, the effect appears to be small, and insufficient to explain the continuous distribution of g-mode pulsators across the HRD found in this and other works (Gaia Collaboration et al., 2023a; Mombarg et al., 2024b; Hey & Aerts, 2024).
It is possible that some of the most strongly tidally affected systems possess induced oscillations that escaped detection. Our methodology is not ideal for the detection of tidally induced pulsations, where the pulsations are perfectly synchronous in frequency with the orbital frequency. Because we rely on harmonic models for the eclipse, such pulsation frequencies are highly vulnerable to being removed during eclipse identification or being confused with an artifact of an imperfectly identified orbital harmonic. With this caveat, even small perturbations from the eclipse orbital frequencies can result in a detectable, extractable oscillation frequency under this methodology (Kemp et al., 2024, 2025).
@@ 3.3 Adjusting Gaia log and log estimates for binarity using eclipse information
Fig. 8 shows the resulting mass vs -mag relations when fixing the central H fraction to different values and when relying on different grids of stellar evolution models. Pre-computed, publicly available spectral synthesis models are used to calculate the -mag based on Gaia’s log and log estimates. The NLTE TLUSTY (Hubeny & Lanz, 2017) spectral synthesis models were used for O-type stars (Lanz & Hubeny, 2003) and B-type stars (Lanz & Hubeny, 2007), while the LTE ATLAS models (Castelli & Kurucz, 2003) were used otherwise. We use the resulting mass-magnitude relations presented in Fig. 8 to test our inferred primary masses by considering that, if our primary masses are correct, a plausible secondary that can reproduce the 2MASS -mag measurement should exist.
@@
Table 4 summarises the number of incompatible targets for each curve in Fig. 8 under all considered primary mass inference schemes. Fig. 9 shows the implied mass ratio distribution for each different value, the M24 and MIST (Choi et al., 2016) stellar tracks, considering four different inferred primary masses obtained: when using the Gaia log and log measurements without modification, when increasing the log measurement by 0.1 dex, when decreasing the log measurements by 0.2 dex, and when using eclipse information to isolate the primary. The region of plausibility, where the mass ratio is between 0 and 1, is highlighted in blue.

### v2

@@ Abstract
In the current age of space-based precision photometry, photometric observations provide the most complete time-domain coverage of binary star systems. Kepler (borucki2010), TESS (ricker2015), and the upcoming PLATO mission (rauer2025) are designed to provide high-precision photometry for large numbers of targets, resulting in the detection and characterisation of tens of thousands of eclipsing binaries (EBs, prsa2011; slawson2011; kirk2016; prsa2022). These EBs allow invaluable constraints to be placed on the binary components and their orbital properties, information that is highly complementary with surface properties obtained through traditional spectroscopic follow-up or, where unavailable or impractical, estimates based on Gaia’s BP/RP low resolution spectra (gaia2023).
@@
A natural synergy between photometric observations of EBs is that with asteroseismology, as the same high-quality time series that allows measurement of eclipse properties also captures flux variations due to pulsations (gaia2023pulsations). At the level of individual stars, this synergy has been proved to be powerful; the combination of EB constraints (hambleton2013; schmid2016; guo2019; sekaran2021; guo2021; aerts2024; kemp2024eclipse; kemp2025_k415) and asteroseismic forward modelling (aerts2018) can provide an invaluable window on fundamental stellar mixing physics (pedersen2021; michielsen2021; michielsen2023; michielsen2024foam). These and similar studies rely heavily on the sensitivity of gravity-mode (g-mode) pulsators to the near-core properties of stars with convective cores and radiative envelopes (main sequence stars with masses greater than 1.2 M⊙). Pressure-mode (p-mode) pulsators (bowman2016; murphy2019; murphy2024) offer complementary information about the outer envelope, making stars in which both g- and p-modes are detected (hybrid pulsators) particularly valuable (kurtz2014; mombarg2020; audenaert2022; kemp2024eclipse; kemp2025_k415).
@@
At the population level, model-independent measurements of the stellar rotation in the near-core region have revealed that main sequence g-mode pulsators are (near-) rigid rotators (aerts2021). Further, using nothing but the dominant g-mode oscillation frequency, an estimate of the stellar mass and near-core rotation rate can be made even from sparse Gaia photometry (hey2024; aerts2025).
@@
Recently, large numbers of eclipsing binaries have been identified and have had their eclipse properties measured using TESS photometry (ijspeert2024_14k) (IJ24). Previous analysis of the asteroseismic potential of this sample was limited to an automated detection procedure based on the number of significant frequencies detected that were not associated with eclipse harmonics. We aim to compile a robust, manually vetted sample of pulsating EBs that is well-characterised in terms of their orbital and stellar properties. Of particular interest is the degree of similarity between the distributions of the subpopulations of pulsating vs non-pulsating EBs, and understanding how binarity can induce bias in the inference of stellar properties. Such a well characterised sample is ideal for applying novel asteroseismic inference methodologies such as the dominant-mode-based estimator for the near-core rotation rate in aerts2025. Further, it provides a starting point for more traditional asteroseismic analyses for targets of high interest and data quality.
@@
In this work, we perform manual vetting of the IJ24 sample in order to identify candidate EB pulsators, and use asteroseismically calibrated stellar models from mombarg2024gaia (M24) to estimate stellar properties for the sample based on their Gaia effective temperature and luminosity estimates. We explore if, and if so how, binarity can result in systematic bias to these results, and confront our mass inferences with 2MASS (skrutskie2006) -magnitude measurements in order to better understand the efficacy of our methodology.
@@ 2.1 Sample definition
Beginning with the 69085-strong eclipsing binary sample of IJ24, we first recover IJ24’s high-quality sample of over 14377111The precise number of stars in our sample differs slightly from IJ24 as we filter both duplicate TESS and Gaia entries, while IJ24 only filters duplicate TESS entries stars. This is achieved by demanding that the automated STAR SHADOW (ijspeert2024) eclipse analyses of IJ24 have completed successfully, filtering for duplicate TESS and Gaia entries, and removing all binaries with that are consistent with 0 eccentricity to within . System lumjinosities are calculated using Gaia DR3 parallax measurements, while other Gaia properties (most notably the effective temperatures) are drawn from esp-hs where available, as it is most suitable for hot stars (7000 K ¡ ¡ 50000 K, gaia2023), and gsp-phot otherwise.
Using the frequency analysis of IJ24, we filtered for stars which have at least one candidate independent frequency of at least 4 mmag. This threshold was designed to enrich our manually vetted sample with pulsators (gaia2023pulsations), as well as improve compatibility with existing empirical asteroseismic methods in the literature (hey2024; aerts2025) relying on dominant g-mode oscillations of at least 4 mmag in amplitude. The 4 mmag limit was originally derived from Gaia DR3 time series (gaia2023pulsations), and so isn’t strictly applicable to TESS photometry. For this reason, we only use this limit for the pre-selection phase resulting in our 4000-strong subsample that we subject to manual vetting; it does not form part of the manual vetting criteria. Most of the scientific discussion in this work will focus on this 4000-target subsample, but note that the stellar property inference was conducted on the entire 14377-target sample of IJ24.
@@
Our 4000-binary subsample of IJ24 was manually inspected in the time and Fourier domains for candidate independent g-mode (751), p-mode (131), or hybrid (48) pulsators. This inspection was done on one year of TESS QLP (huang2020I; huang2020II; kunimoto2021) or SPOC (kunimoto2021) data per target, where the data used was selected based on the following priority ranking from highest to lowest:
@@ 2.2 Stellar parameter estimation
Using conditional normalising flows (CNFs, winkler2019; hon2024) trained on the M24 stellar evolution models, we estimate masses (), convective core masses (), radii (), and central H fractions () for the 12508 targets with Gaia effective temperatures and luminosities. This estimation is based on the previously described Gaia log and log estimates, and is performed for the three metallicities () available, marginalising over the initial (rigid) rotations () and values for the exponential overshoot () varied in the M24 grids.
@@
where is the height above the convective boundary, is the pressure scale height at the convective boundary, and is the diffusivity coefficient at the edge of the convective boundary (freytag1996; herwig2000; mombarg2024). Following M24, we uniformly resample the evolution models in to avoid bias towards the zero-age main sequence (ZAMS) and terminal-age main sequence (TAMS) caused by the numerics of the evolution models. For all targets, a kernel density estimate (KDE) is then performed to obtain the maximum likelihood estimate and its uncertainty by uniformly distributing 800 points in each stellar parameter (, , , log, , and ).
@@ 3.1 Basic stellar and orbital properties
The distribution of our stars across the Hertzsprung-Russel diagram (HRD) is shown in Fig. 1. The upper left panel also shows solar metallicity (=0.014) evolutionary tracks from 1.2 ⊙ to 9 M⊙ from M24. P-values for 2-sample Kolmogorov-Smirnov (KS) tests (massey1951kolmogorov; hodges1958) applied to all subsample pairings can be found in Table 2. Unless otherwise stated, the p-values discussed in the main text always adopt the null hypothesis that the relevant pairing of subsamples is drawn from the same underlying distribution; hence, a high p-value implies a high degree of similarity between the distributions, while a low value indicates significant discrepancies. The lower panels highlight the distribution of the g-mode, p-mode, and hybrid pulsators.
@@
The distribution of the g-mode pulsators is relatively continuous across the HRD; while the 2-sample KS test probability with the non-pulsator class is low (0.045 for log), it is also significantly higher than that of either the p-mode or hybrid samples (0.006 and 0.001, respectively). Qualitatively, there is a slight bimodality in log for the g-mode stars that is absent for the non-pulsators. This may be the signature of the Dor (guzik2000; dupret2005; uytterhoeven2011; bouabid2013) and slowly-pulsating B-star (SPB, waelkens1991; dziembowski1993; gautschy1993; miglio2007) instability regions. However, we still must conclude that the distribution of g-mode pulsators across the HRD appears more similar to the distribution of non-pulsators than we would expect for a (pure) sample of g-mode pulsators confined to the traditional Dor, SPB, and Cep (dziembowski1993bcep; stankov2005) instability regions. Pulsators outside these regions have previously been found by combining Gaia astrometry with Kepler or TESS photometry (gaia2023; balona2023; hey2024; fritzewski2024ubc1; mombarg2024). It is possbile that this blurring of the traditional instability regions could be due to fundamental limitations to the precision and accuracy of Gaia estimates in the SPB regime. However, it is at least as plausible that this phenomena is due to the physical limitations of our excitation models, which are missing proper treatments for physics such as stellar rotation and tides aerts2024.
In the context of our eclipsing binary sample, a key question is whether tidal interactions (fuller2012; fuller2017; guo2021; aerts2024) are a dominant force for the g-mode pulsator classes becoming blurred. Alternative explanations include non-binary related missing physics in mode excitation models, such as rotation, but also the possibility of a relatively high level of contamination by rotational variables, which, even with manual vetting, can be difficult to distinguish from genuine g-mode pulsators in TESS photometry (audenaert2021; hey2024).
@@
The period-eccentricity distribution of the 4000-target vetted sample is shown in Fig. 2. G-mode pulsators, including hybrids, are preferentially found at short orbital periods compared to both the non-pulsator subclass and the p-mode pulsators. This contrasts with Fig. 1, which showed that in terms of stellar surface properties, the distribution of the hybrids and p-mode were far more similar than that of hybrids and g-modes. The p-mode pulsators peak at longer orbital periods and generally have a flatter distribution that is more consistent with the non-pulsator subclass (p-value = )) than either the g-mode (p-value = ) or hybrid pulsators (p-value = )). This preference for short-period, circular binaries is consistent with increased tidal efficiency in the presence of g-mode pulsations, which would lead to more rapid spin-orbit coupling and circularisation (rogers2013; ogilvie2014).
@@
Comparing the upper and lower panels of Fig. 3, there does not appear to be a strong correlation between tidal variability and the g-mode pulsators. This is not consistent with tidal effects dominantly affecting the excitation of g-modes in our sample. If this were the case, we would expect more g-mode pulsators in systems where tidal variability is the strongest. Instead, the g-mode pulsators are distributed relatively evenly between low and moderate values (0-0.4) of our tidal morphology parameter. Comparing the different pulsator classes (see Fig. 11), we note that the g-mode pulsators prefer higher values of the tidal-morphology parameter than the p-mode pulsators (p-value = 0.004), while the hybrids closely match the g-mode distribution (p-value = 0.73, compared to a p-mode — hybrid p-value of 0.03). While this does constitute evidence for g-mode pulsators being affected by tides more than targets without g-mode pulsations, the effect appears to be small, and insufficient to explain the continuous distribution of g-mode pulsators across the HRD found in this and other works (gaia2023pulsations; mombarg2024gaia; hey2024).
It is possible that some of the most strongly tidally affected systems possess induced oscillations that escaped detection. Our methodology is not ideal for the detection of tidally induced pulsations, where the pulsations are perfectly synchronous in frequency with the orbital frequency. Because we rely on harmonic models for the eclipse, such pulsation frequencies are highly vulnerable to being removed during eclipse identification or being confused with an artifact of an imperfectly identified orbital harmonic. With this caveat, even small perturbations from the eclipse orbital frequencies can result in a detectable, extractable oscillation frequency under this methodology (kemp2024eclipse; kemp2025_k415).
@@ 3.3 Adjusting Gaia log and log estimates for binarity using eclipse information
Fig. 8 shows the resulting mass vs -mag relations when fixing the central H fraction to different values and when relying on different grids of stellar evolution models. Pre-computed, publicly available spectral synthesis models are used to calculate the -mag based on Gaia’s log and log estimates. The NLTE TLUSTY (hubeny2017) spectral synthesis models were used for O-type stars (lanz2003) and B-type stars (lanz2007), while the LTE ATLAS models (castelli2003) were used otherwise. We use the resulting mass-magnitude relations presented in Fig. 8 to test our inferred primary masses by considering that, if our primary masses are correct, a plausible secondary that can reproduce the 2MASS -mag measurement should exist.
@@
Table 4 summarises the number of incompatible targets for each curve in Fig. 8 under all considered primary mass inference schemes. Fig. 9 shows the implied mass ratio distribution for each different value, the M24 and MIST (choi2016) stellar tracks, considering four different inferred primary masses obtained: when using the Gaia log and log measurements without modification, when increasing the log measurement by 0.1 dex, when decreasing the log measurements by 0.2 dex, and when using eclipse information to isolate the primary. The region of plausibility, where the mass ratio is between 0 and 1, is highlighted in blue.